How many moles of silver nitrate are needed to add to 35 moles of potassium phosphate to make the final products


1
Expert's answer
2021-04-19T05:07:26-0400

K3PO4 + 3AgNO3=3KNO3+Ag3PO4

let's make a proportion for finding moles of silver nitrate, take it for x to moles of silver nitrate

"\frac{x}{1}=\frac{35}{3}"

x=11.67moles of silver nitrate